# Text to Band Gap: Pre-trained Language Models as Encoders for Semiconductor Band Gap Prediction

### v2

@@ Abstract
We investigate the use of transformer-based language models, RoBERTa, T5, and LLaMA, for predicting the band gaps of semiconductor materials directly from textual representations that encode key material features such as chemical composition, crystal system, space group, number of atoms per unit cell, valence electron count, and other relevant electronic and structural properties. Quantum chemistry simulations such as Density Functional Theory (DFT) provide accurate predictions but are computationally intensive, limiting their feasibility for large-scale materials screening. Shallow machine learning (ML) models offer faster alternatives but typically require extensive data preprocessing to convert non-numerical material features into structured numerical inputs, often at the cost of losing critical descriptive information. In contrast, our approach leverages pretrained language models to process textual data directly, eliminating the need for manual feature engineering. We construct material descriptions in two formats: structured strings that combine key features in a consistent template, and natural language narratives generated using the ChatGPT API. For each model, we append a custom regression head and perform task-specific finetuning on a curated dataset of inorganic compounds. Our results show that finetuned language models, particularly the decoder-only LLaMA-3 architecture, can outperform conventional approaches in prediction accuracy and flexibility, achieving a mean absolute error (MAE) of 0.248 eV, root mean squared error (RMSE) of 0.345 eV, and $R^{2}$ of 0.891, compared to the best shallow ML baseline (XGBoost), which achieved an MAE of 0.318 eV, RMSE of 0.537 eV, and $R^{2}$ of 0.838. Notably, LLaMA-3 achieves competitive accuracy with minimal fine-tuning, suggesting its architecture enables more transferable representations for scientific tasks. This work demonstrates the effectiveness of finetuned language models for scientific property prediction and provides a scalable, language-native framework for materials informatics.
@@ 1 Introduction
Machine learning (ML) methods have become powerful tools for addressing the computational challenges of DFT. Shallow ML models, such as Random Forest and Support Vector Regression, are commonly used to predict materials properties like band gaps based on material descriptors 7, 8, 9. These models offer a cost-effective alternative to solving the full quantum mechanical equations, significantly reducing computational overhead. However, these models often struggle with non-numerical features, requiring extensive preprocessing to convert material properties into numerical formats. This reliance on extensive preprocessing and feature engineering not only adds complexity but also risks discarding nuanced or qualitative information, such as symmetry, bonding environments, or textual metadata, that could be valuable for accurate property prediction 10, 11. Recent advances in deep learning, particularly Graph Neural Networks (GNNs) 12, have improved the ability to capture the interconnectivity of atomic structures by representing them as graphs. However, this approach requires the conversion of atomic arrangements into graph representations, which adds another layer of preprocessing. Furthermore, GNNs still face limitations in integrating non-numerical properties, such as compound names, into the training process without additional preprocessing steps. These challenges underscore the need for approaches that can seamlessly handle both numerical and non-numerical features in material property predictions while minimizing complex preprocessing requirements.
@@
In this study, we explore the use of transformer-based language models, RoBERTa, T5, and LLaMA-3, to predict the band gaps of semiconductor materials directly from textual descriptions. These models enable the direct encoding of structured or natural language representations of materials, such as chemical composition, crystal symmetry, and electronic features, without requiring conventional feature engineering. While pretrained language models possess strong linguistic priors, we emphasize that fine-tuning on domain-specific objectives is essential for adapting them to materials property prediction. We implement task-specific regression heads on top of each model and finetune them on a curated dataset of inorganic compounds. This approach allows models to learn mappings from text-based input to scalar band gap values. These models provide a flexible and generalizable framework for property prediction from textual materials data, extending the application of language models beyond conventional natural language tasks into scientific domains such as materials informatics.
@@ 2.1 RoBERTa
RoBERTa (Robustly Optimized BERT Approach) is a transformer-based encoder-only language model that builds upon the BERT architecture by refining pretraining strategies and removing restrictive design choices 20. It is widely recognized for its ability to capture rich contextual relationships between tokens through deep bidirectional self-attention, making it highly effective for a range of natural language processing tasks.
RoBERTa is pretrained using the masked language modeling (MLM) objective, where a subset of tokens is masked and the model learns to predict them based on surrounding context. Unlike BERT, RoBERTa omits the next-sentence prediction task and instead trains on longer, contiguous sequences of text 20. It also introduces dynamic masking, larger batch sizes, and improved sentence packing strategies, which collectively enhance the generality and robustness of the learned representations. Tokenization is performed using a byte-level Byte-Pair Encoding (BPE) tokenizer, which allows the model to effectively handle multilingual text, rare subwords, and non-standard Unicode characters.
In this study, we leverage the RoBERTa-base variant, comprising 12 transformer layers, 12 attention heads per layer, and a hidden dimension of 768. While RoBERTa provides strong pretrained representations, our framework emphasizes the necessity of objective-specific fine-tuning to adapt the model to domain-specific tasks such as band gap prediction. The model’s encoder-only structure offers advantages in processing structured text inputs efficiently, particularly in cases where outputs are scalar values rather than generated sequences.
@@
T5 (Text-to-Text Transfer Transformer) is a unified encoder-decoder architecture that reframes all NLP tasks—including classification, translation, and regression—as text-to-text problems 21. In this study, we utilize the T5-small variant, which consists of 6 encoder and 6 decoder layers, each with a hidden size of 512 and 8 attention heads. Unlike encoder-only models such as RoBERTa, T5 processes input and output sequences through separate but coordinated transformer stacks, allowing for flexible sequence-to-sequence modeling and richer bidirectional context capture within the encoder.
T5 is pretrained using a span corruption objective, in which contiguous spans of tokens in the input are replaced by unique sentinel tokens, and the model learns to reconstruct the missing content. This span-level objective fosters stronger global reasoning compared to token-level masking and supports robust performance on tasks requiring understanding of longer context windows. Architectural enhancements include the use of relative positional embeddings rather than absolute positions, parameter sharing between input and output embeddings, and simplified layer normalization without additive bias. These design choices collectively reduce model complexity and improve generalization.
Tokenization in T5 is performed using SentencePiece, a subword tokenizer that supports language-agnostic processing and eliminates the need for whitespace-based token segmentation 22, 23. It is particularly useful for scientific or symbolic inputs, as it can encode rare tokens and domain-specific terms with high fidelity. During training, subword regularization introduces stochasticity in tokenization, enhancing robustness by exposing the model to a diverse range of token combinations.
While T5’s encoder-decoder structure is typically leveraged for generative tasks, it is also adaptable for scalar prediction tasks like band gap regression. In our implementation, we repurpose the encoder to process material descriptions and append a custom regression head to the encoder output.
@@ 2.3 LLAMA-3
LLaMA-3.2-1B is a lightweight, decoder-only transformer model derived from Meta’s LLaMA-3 family of large language models 24, 25. Designed to provide a strong performance-efficiency trade-off, this version incorporates approximately 1 billion parameters and benefits from structured pruning and knowledge distillation from larger models. The architecture emphasizes both representational depth and inference efficiency, making it suitable for scalable applications beyond conventional text generation tasks.
Unlike encoder-only models like RoBERTa or encoder-decoder frameworks such as T5, LLAMA-3 operates in an autoregressive, decoder-only configuration. Despite its unidirectional nature, LLAMA-3 can learn rich representations through deeply stacked attention layers and efficient architectural enhancements. The model employs several innovations that distinguish it from earlier transformers: it uses RMSNorm for pre-normalization to improve training stability, SwiGLU as a more expressive activation function, and Rotary Positional Embeddings (RoPE) to better generalize to longer sequences 26. In some variants, Grouped Query Attention (GQA) is used to reduce computational overhead while maintaining strong modeling capacity.
LLaMA-3 tokenization is performed using BPE via SentencePiece 22, which enables robust handling of diverse character sets and subword units. Numerals are decomposed into individual digit tokens, and rare or unknown characters are represented using byte-level encodings. These choices ensure flexibility in modeling scientific texts that may include specialized symbols, units, or chemical formulas.
Although LLAMA-3 was originally developed for autoregressive text generation, its modular architecture and robust token-level representation capabilities make it a versatile candidate for regression tasks. The model is particularly well-suited for structured and semi-structured textual input, where consistent formatting allows it to encode domain-relevant relationships. In our context, a custom regression head is appended to the decoder’s final hidden states to adapt the model for scalar band gap prediction. As with other language models in this study, fine-tuning on task-specific objectives is necessary to align the model’s pretrained linguistic knowledge with the specialized needs of materials property prediction.
@@ 2.4 Dataset
In this study, we utilized the AFLOW database, a comprehensive open repository for computational materials science that contains extensive information on inorganic crystalline materials and their properties 27, 28. Band gap calculations in AFLOW combine first-principles methods with empirical corrections through an automated workflow. The framework uses VASP to perform DFT calculations with the GGA-PBE functional for standard compounds while applying the GGA+U method for strongly correlated systems containing d- and f-block elements. To address GGA’s tendency to underestimate band gaps, AFLOW employs an empirical correction scheme based on a linear regression model derived from 102 benchmark compounds with known experimental values29, 27, 30. This systematic approach, along with the database’s vast size and rich feature space, makes AFLOW particularly well-suited for ML tasks aimed at material property prediction.
For our specific analysis, we selected a subset of 27,600 materials with band gap values ranging between 0 and 5 eV (inclusive). This range was chosen because it encompasses the most relevant band gap values for semiconductors, which are of particular interest in materials science and electronic applications. The lower bound of 0 eV represents materials with metallic behavior, where there is no electronic band gap. Materials that have band gap higher than 5 eV are insulating materials, which have too large band gaps and do not conduct electricity under normal conditions 31, 3. By focusing on materials within this range of 0-5 eV band gaps, we ensure that our model targets materials with practical applications in electronics and optoelectronics.
@@ 2.5 Text Data Format
Both formats were applied uniformly across all three models: RoBERTa, T5, and LLaMA-3. The models processed these inputs through their native tokenization pipelines, without additional handcrafted feature engineering. For each model, we appended a custom regression head to enable scalar prediction of band gap values and performed fine-tuning on the downstream task using the corresponding textual inputs.
@@ 2.6 Input Features
The chemical formula represents the basic building blocks of the material, providing critical information about its stoichiometry and composition. The nature and type of atoms constituting the material greatly influence its electronic properties32, 9, 33. For example, the number of valence electrons of each species is crucial for band gap predictions34. The total number of valence electrons per unit cell helps predict how these electrons will contribute to the band structure, which is directly related to the band gap.
Additionally, we paid particular attention to structural features, including crystal class, family, and system, as well as lattice parameters, which include the dimensions and angles of the unit cell. These factors not only shape the arrangement of atoms and their interactions but also define the symmetry and geometric properties of the crystal, directly influencing the distribution of electronic states within the energy bands 35, 36, 30. The space group and point group information were included to account for the effects of symmetry on electronic states and band splitting. Magnetic properties, such as atom magnetic moments and cell magnetization, were also considered due to their relationship with spin distribution and electronegativity3. We specifically chose properties derived from relaxed structures in which the structural configurations have been optimized to minimize energy and stress, ensuring that the atoms are in their equilibrium positions.
@@ 3.1 Framework
Figure 1: Overview of the proposed band gap prediction framework. a The pipeline starts from the AFLOW dataset, followed by feature selection, dataset preparation, and LLM model training for final band gap prediction. b Two input formats are illustrated. string-based representation using direct feature values and description-based format generated by GPT-3.5 turbo. c Visualization of the finetuning process. The input text undergoes tokenization and embedding through a pretrained LLM, followed by a custom regression head for prediction. d Demonstrates the Transformer encoder and the multi-head attention mechanism with Query (Q), Key (K), and Value (V) operations.
We developed a language model-based framework to predict the band gaps of semiconductor materials using transformer-based language models: RoBERTa, T5, and LLaMA-3. As shown in Figure 1a, we constructed the dataset from the AFLOW database by extracting relevant material features and transforming them into textual formats suitable for language model inputs. Two types of textual representations were used (Figure 1b): a structured string format that followed a fixed template, and a more flexible natural language description generated using the GPT-3.5 Turbo API. These formats enabled us to assess how each model handles both highly regular and semantically rich input styles. Additional examples of both the structured strings and natural language descriptions are provided in the Supporting Information.s
At the core of our approach are transformer-based language models, each differing in architecture and tokenization strategy. While RoBERTa uses an encoder-only structure, T5 adopts an encoder-decoder design, and LLaMA-3 operates with a decoder-only architecture. Each model tokenizes the input text using its native tokenizer, such as byte-level BPE or SentencePiece, and processes the sequence to generate contextual embeddings (Figure 1c). These embeddings are passed through a custom regression head to produce a scalar band gap prediction, as shown in Figure 1d. The design of this framework allows for direct comparison across model types and input formats while minimizing the need for handcrafted features or domain-specific encodings.

### v3

@@ Abstract
We investigate transformer-based language models, including RoBERTa, T5, Llama-3, and MatSciBERT, for predicting the band gaps of semiconductor materials directly from textual descriptions. The inputs encode key material features, such as chemical composition, crystal system, space group, and other structural and electronic properties. Unlike shallow machine learning models, which require extensive feature engineering, or Graph Neural Networks, which rely on graph representations derived from atomic coordinates, pretrained language models can process textual inputs directly, eliminating the need for manual feature preprocessing or structure-based encoding. Material descriptions were constructed in two formats: structured strings with a consistent template and natural language narratives generated via the ChatGPT API. Each model was augmented with a custom regression head and finetuned for band gap prediction task. Language models of different architectures and parameter sizes were all able to predict band gaps from human-readable text with strong accuracy, achieving MAEs in the range of 0.25–0.33 eV, highlighting the success of this approach for scientific regression tasks. Finetuned Llama-3, with 1.2 billion parameters, achieved the highest accuracy (MAE 0.248 eV, $R^{2}$ 0.891). MatSciBERT, pretrained on materials science literature, reached comparable performance (MAE 0.288 eV, $R^{2}$ 0.871) with significantly fewer parameters (110 million), emphasizing the importance of domain-specific pretraining. Attention analysis shows that both models selectively focus on compositional and spin-related features while de-emphasizing geometric features, reflecting the difficulty of capturing spatial information from text. These results establish that pretrained language models can effectively extract complex feature-property relationships from textual material descriptions, providing a scalable, language-native framework for materials informatics.
@@ 1 Introduction
Machine learning (ML) methods have become powerful tools for addressing the computational challenges of DFT. Shallow ML models, such as Random Forest and Support Vector Regression, are commonly used to predict materials properties like band gaps based on material descriptors 7, 8, 9. These models offer a cost-effective alternative to solving the full quantum mechanical equations, significantly reducing computational overhead. However, these models often struggle with non-numerical features, requiring extensive preprocessing to convert material properties into numerical formats. This reliance on extensive preprocessing and feature engineering not only adds complexity but also risks discarding nuanced or qualitative information, such as symmetry, bonding environments, or textual metadata, that could be valuable for accurate property prediction 10, 11. Recent advances in deep learning, particularly Graph Neural Networks (GNNs) 12, have significantly enhanced the modeling of structure–property relationships by representing materials as atomic graphs that encode interatomic connectivity. However, these methods inherently rely on explicit structural information and require the conversion of atomic configurations into graph representations, introducing additional preprocessing steps and limiting their applicability to systems with well-defined crystal structures. Furthermore, GNNs still face limitations in integrating non-numerical properties, such as compound names, into the training process without additional preprocessing steps. These challenges underscore the need for approaches that can seamlessly handle both numerical and non-numerical features in material property predictions while minimizing complex preprocessing requirements.
@@
In this study, we explore the use of transformer-based language models, RoBERTa, T5, Llama-3, and MatSciBERT to predict the band gaps of semiconductor materials directly from textual descriptions. These models enable the direct encoding of structured or natural language representations of materials, such as chemical composition, crystal symmetry, and electronic features, without requiring conventional feature engineering. While pretrained language models possess strong linguistic priors, we emphasize that fine-tuning on domain-specific objectives is essential for adapting them to materials property prediction. We implement task-specific regression heads on top of each model and finetune them on a curated dataset of inorganic compounds. This approach allows models to learn mappings from text-based input to scalar band gap values. These models provide a flexible framework for property prediction from textual materials data, extending the application of language models beyond conventional natural language tasks into scientific domains such as materials informatics.
@@ 2.1 RoBERTa
RoBERTa (Robustly Optimized BERT Approach) is an encoder-only transformer that refines BERT’s pretraining strategy through dynamic masking and longer sequence training 20. In this work, we employ the RoBERTa model (12 layers, 12 attention heads, 768 hidden units) as one of the backbone architectures for band gap prediction. Its bidirectional attention enables effective contextual encoding of structured text inputs derived from material features. Rather than emphasizing linguistic tasks, our framework fine-tunes RoBERTa to learn correlations between textual feature descriptions and scalar band gap values. This objective-specific adaptation allows the model to repurpose general language representations for scientific property prediction.
@@
T5 (Text-to-Text Transfer Transformer) is a unified encoder–decoder architecture that reformulates all NLP tasks as text-to-text transformations 21. We employ the T5-small (6 encoder and 6 decoder layers, 512 hidden size, 8 attention heads) to explore how a sequence-to-sequence framework can be applied to scientific regression problems. Throughout this manuscript, we refer to this model simply as T5.
Pretrained with a span corruption objective, T5 learns to reconstruct masked text segments, promoting stronger global reasoning than token-level masking. In our framework, we utilize only the encoder output and append a custom regression head for band gap prediction. This setup leverages T5’s contextual encoding ability while avoiding unnecessary generative complexity, allowing efficient adaptation to structured material descriptions and scalar property prediction.
@@ 2.3 Llama-3
Llama-3 is a family of decoder-only transformer models developed by Meta, designed to provide state-of-the-art performance with efficient scaling across parameter sizes 22, 23. In this work, we specifically use the lightweight Llama-3.2-1B variant (approximately one billion parameters, embedding size 2048), which offers a strong balance between performance and computational efficiency. Throughout this manuscript, we refer to this model simply as Llama-3.
We adapt Llama-3 by attaching a custom regression head to the decoder output for band gap prediction. Architectural components such as RMSNorm, SwiGLU activation, and Rotary Positional Embeddings improve training stability and contextual reasoning, while the SentencePiece-based tokenizer effectively encodes scientific symbols and numerical expressions. This configuration enables Llama-3 to extend its pretrained linguistic representations toward modeling structured material descriptions for quantitative property prediction.
2.4 MatSciBERT
MatSciBERT is a BERT-based language model pretrained on materials science literature to capture domain-specific terminology, chemical formulas, and structural descriptions 24. It uses the standard BERT-base architecture with 12 encoder layers, 768-dimensional embeddings, 12 attention heads, and a 3072-dimensional feed-forward layer, while its masked language modeling pretraining is optimized for the symbolic and long-tail vocabulary common in scientific texts. This provides stronger representations for materials-related language than general-purpose models.
In our framework, we employ the MatSciBERT encoder with a custom regression head to predict band gaps from structured material descriptions. This design leverages MatSciBERT’s domain-aware embeddings while maintaining architectural simplicity, allowing efficient adaptation to property prediction tasks within materials informatics.
2.5 Shallow ML Models
We implemented three conventional regression algorithms, Random Forest (RF), Support Vector Regression (SVR), and Extreme Gradient Boosting (XGBoost), using the same dataset of structured material descriptors. All models were trained to predict the band gap values from numerical features extracted from the materials database.
The RF model was implemented using scikit-learn’s RandomForestRegressor. The model constructs an ensemble of decision trees trained on randomly sampled subsets of both data and features, and outputs the average prediction across all trees. A five-fold cross-validated grid search was used to optimize key hyperparameters. The final configuration employed 1,000 trees, a maximum depth of 50, a minimum of two samples per split, and two samples per leaf.
For the SVR, we used scikit-learn’s SVR implementation with a radial basis function (RBF) kernel. The model projects the input features into a high-dimensional space where a linear regression is performed within an $\epsilon$-insensitive margin. Grid search with five-fold cross-validation determined the optimal hyperparameters: a penalty parameter $C=5000$, $\epsilon=0.1$, and an RBF kernel. The high value of $C$ enables the model to fit complex nonlinear patterns, while $\epsilon$ controls the tolerance for prediction deviations near the regression boundary.
The XGBoost model was implemented using the XGBRegressor from the xgboost library. XGBoost builds an ensemble of boosted decision trees, where each successive tree is trained to correct the residual errors of the previous ensemble. Grid search with five-fold cross-validation identified the best configuration: 2,000 trees, maximum depth of 9, learning rate of 0.1, subsample ratio of 1.0, and column sampling ratio of 0.6. These settings yield deep, expressive trees while maintaining regularization through feature subsampling. The model objective was set to minimize the squared error loss.
@@ 2.6 Dataset
In this study, we utilized the AFLOW database, a comprehensive open repository for computational materials science that contains extensive information on inorganic crystalline materials and their properties 25, 26. Band gap calculations in AFLOW combine first-principles methods with empirical corrections through an automated workflow. The framework uses VASP to perform DFT calculations with the GGA-PBE functional for standard compounds while applying the GGA+U method for strongly correlated systems containing d- and f-block elements. To address GGA’s tendency to underestimate band gaps, AFLOW employs an empirical correction scheme based on a linear regression model derived from 102 benchmark compounds with known experimental values27, 25, 28. This systematic approach, along with the database’s vast size and rich feature space, makes AFLOW particularly well-suited for ML tasks aimed at material property prediction.
For our specific analysis, we selected a subset of 27,600 materials with band gap values ranging between 0 and 5 eV (inclusive). This range was chosen because it encompasses the most relevant band gap values for semiconductors, which are of particular interest in materials science and electronic applications. The lower bound of 0 eV represents materials with metallic behavior, where there is no electronic band gap. Materials that have band gap higher than 5 eV are insulating materials, which have too large band gaps and do not conduct electricity under normal conditions 29, 3. By focusing on materials within this range of 0-5 eV band gaps, we ensure that our model targets materials with practical applications in electronics and optoelectronics.
@@ 2.7 Text Data Format
Both formats were applied uniformly across all three models: RoBERTa, T5, Llama-3, and MatSciBERT. The models processed these inputs through their native tokenization pipelines, without additional handcrafted feature engineering.
@@ 2.8 Input Features
The chemical formula represents the basic building blocks of the material, providing critical information about its stoichiometry and composition. The nature and type of atoms constituting the material greatly influence its electronic properties30, 9, 31. Specifically, the elemental types and their ratios determine atomic energy levels, such as s, p, and d orbitals, as well as bond types like covalent and ionic bonds, and electronegativity differences. These factors affect the relative positions and separation of the valence and conduction bands. In ionic compounds, a larger electronegativity difference leads to a greater energy separation between the valence band, which is primarily formed by anion orbitals, and the conduction band, which is mainly formed by cation orbitals. This typically results in a larger band gap. For example, II-VI compounds often have larger band gaps compared to III-V compounds32. For example, the number of valence electrons of each species is crucial for band gap predictions33. The total number of valence electrons per unit cell governs how the electronic bands are filled and directly influences the position of the Fermi level34. Differences in valence electron count and orbital configuration among elements or compounds can lead to significant variations in both the magnitude and nature of the band gap. Materials with stronger orbital overlap and higher structural symmetry tend to have wider bands and smaller, often direct, band gaps, whereas weaker overlap or lower symmetry can result in indirect or larger gaps, ultimately determining the optical and transport properties of the material35.
Additionally, we paid particular attention to structural features, including crystal class, family, and system, as well as lattice parameters, which include the dimensions and angles of the unit cell. These factors not only shape the arrangement of atoms and their interactions but also define the symmetry and geometric properties of the crystal, directly influencing the distribution of electronic states within the energy bands 36, 37, 28. The crystal’s symmetry determines the shape of the Brillouin zone and band degeneracy. High symmetry, as seen in cubic systems, often leads to high band degeneracy at high-symmetry points, where multiple electrons have the same energy, resulting in a relatively simple band structure. In contrast, lower symmetry, as in orthorhombic systems, can cause band splitting, which may change the band gap’s size and type38. The space group and point group information were included to account for the effects of symmetry on electronic states and band splitting. Magnetic properties, such as atom magnetic moments and cell magnetization, were also considered due to their relationship with spin distribution and electronegativity3. We specifically chose properties derived from relaxed structures in which the structural configurations have been optimized to minimize energy and stress, ensuring that the atoms are in their equilibrium positions.
@@ 3.1 Framework
Figure 1: Overview of the proposed band gap prediction framework. a The pipeline starts from the AFLOW dataset, followed by feature selection, dataset preparation, and LLM model training for final band gap prediction. b Two input formats are illustrated. string-based representation using direct feature values and description-based format generated by GPT-3.5 turbo. c Visualization of the finetuning process. The input text undergoes tokenization and embedding through multiple model architectures (RoBERTa, T5, Llama-3, MatSciBERT), followed by a custom regression head for prediction. d Demonstrates the Transformer encoder and the multi-head attention mechanism with Query (Q), Key (K), and Value (V) operations.
We developed a language model-based framework to predict the band gaps of semiconductor materials using transformer-based language models: RoBERTa, T5, Llama-3, and MatSciBERT. As shown in Figure 1a, we constructed the dataset from the AFLOW database by extracting relevant material features and transforming them into textual formats suitable for language model inputs. Two types of textual representations were used (Figure 1b): a structured string format that followed a fixed template, and a more flexible natural language description generated using the GPT-3.5 Turbo API. These formats enabled us to assess how each model handles both highly regular and semantically rich input styles. Additional examples of both the structured strings and natural language descriptions are provided in the Supporting Information.
At the core of our approach are transformer-based language models, each differing in architecture, tokenization strategy, and training strategy. Our approach is built on transformer-based language models with distinct architectures and training schemes. RoBERTa and MatSciBERT share an encoder-only design, but MatSciBERT is additionally pretrained on materials science texts for domain-specific understanding. T5 uses an encoder–decoder structure, while Llama-3 is a decoder-only model. Each model tokenizes the input text using its native tokenizer, such as byte-level BPE or SentencePiece, and processes the sequence to generate contextual embeddings (Figure 1c). These embeddings are passed through a custom regression head to produce a scalar band gap prediction.